The principal is 5,700
Term is 60 days
Interest $95
What would the interest rate be.

Answer by Boreal(15235)

Using 360 day year, which some do.
I=prt
95=5700*r*(1/6)=950r
r=95/950=10%
==========
Using a 365 day year, 95=5700*60/365)*r
r=95/936.99
=10.14%
given the 10% answer, the problem probably wants a 360 day year.